OpenStudy (gabylovesyou):

@sleepyjess The function c(x) = 3x − 4 determines how many cupcakes need to be purchased for a birthday party, where x is the number of kids at the party. The party planner uses d(c(x)) to find the amount of money to bring for the cupcake purchase. The function d(x) = 2x + 5. Solve for how much money to bring when there are 7 kids at the party.

OpenStudy (sleepyjess):

First, do you know what d(c(x)) means?

OpenStudy (here_to_help15):

We have two individual functions.

OpenStudy (here_to_help15):

c(x)=3x-4, how many cupcakes to bring.

OpenStudy (gabylovesyou):

yes it means d(3*7 - 4) = 2*(3*7 - 4) + 5

OpenStudy (sleepyjess):

Yep :(

OpenStudy (sleepyjess):

*:)

OpenStudy (gabylovesyou):

but i keep solving it wrong ._.

OpenStudy (here_to_help15):

d(x)=2x+5, how much money for a certain number of cupcakes.

OpenStudy (sleepyjess):

Ok, your substituting looks good.

OpenStudy (sleepyjess):

So do you know where you keep going wrong?

OpenStudy (gabylovesyou):

no idea... first i start with the parenthesis right ?

OpenStudy (here_to_help15):

To determine how much money the party planner needs means that they use the function d(c(x)) like @sleepyjess said

OpenStudy (sleepyjess):

Yep, start with 3*7 in the parentheses

OpenStudy (gabylovesyou):

Nvm i got it lol its 39

OpenStudy (here_to_help15):

Which is,d(3*7 - 4) = 2*(3*7 - 4) + 5 as you said

OpenStudy (here_to_help15):

Sorry if i wasnt much help hehe cya :P :)

OpenStudy (gabylovesyou):

Thank you both :)

OpenStudy (sleepyjess):

Yep, 39 is correct :) No problem!

OpenStudy (here_to_help15):

Np : )

OpenStudy (here_to_help15):

hehe ill take ur thank you as my medal :)

OpenStudy (sleepyjess):

Wow this was my 2000th question answered!